OpenStudy (anonymous):

What is the probability that a fair coin will show “heads” 2 out of 4 times?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

the answer is 3/8

OpenStudy (anonymous):

i know that there is a possible of 16

OpenStudy (adilalvi):

0.5

OpenStudy (anonymous):

2^4 is 16 possible then there is 6 ways that you can get head

OpenStudy (kropot72):

The binomial distribution can solve this: \[P(2\ heads)=\left(\begin{matrix}4 \\ 2\end{matrix}\right)0.5^{2}\times 0.5^{2}=\frac{6}{2^{4}}\]